OpenStudy (anonymous):

Need help don't understand this kind of stuff really well.... There are ____ mol in 50g of CaCO3

OpenStudy (aaronq):

Assuming that you know what a mole is, use the formula: \(\sf \large moles=\dfrac{mass ~(in~g)}{Molar~mass ~(in~g/mol)}\) so you have to first find the molar mass of CaCO3 with the periodic table.

OpenStudy (anonymous):

That's wat I'm lost about

OpenStudy (aaronq):

for the molar mass of a compound, you add the molar of each individual atom in proportion to the number of atoms present. So for example, water, \(H_2O\) has 2 H and 1 O Molar mass of water= 2*(1 g/mol) + 1*(16 g/mol)= 18 g/mol

OpenStudy (aaronq):

For your compound you have, 1 Ca, 1 C, and 3 O

OpenStudy (anonymous):

Okay so what would the problem look like because I have to do 5 more like this.. And just what to say is thanks for your help

OpenStudy (aaronq):

so you use the molar masses listed on the periodic table \(\sf moles=\dfrac{50~g} {\underbrace{(40~g/mol)}_{Ca}+\underbrace{(12~g/mol)}_{C}+\underbrace{(3*18~g/mol)}_{O} }\)

OpenStudy (aaronq):

oh wait O is 16 g/mol not 18 g/mol

OpenStudy (anonymous):

Thanks because you explain it better than my teacher did and is is 61.25??

OpenStudy (aaronq):

no problem (sorry for the late reply, i'm not receiving notifications) and the answer would be much less than 61.25 \(\sf moles=\dfrac{50~g}{100~g/mol}=0.5~mol\)